OpenStudy (anonymous):

Frank paid $22.40 for a sweatshirt. It had been discounted by 30%. What was the regular price of the sweatshirt? Solve to build an equation to describe the situation.

OpenStudy (anonymous):

22.40 is 70% of the original price thus 70%x=22.40 or 0.7x = 22.40 solve we have x = 32

OpenStudy (anonymous):

Yep. I only needed an equation, you did it a little backwards.

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

lol thats what i thought

OpenStudy (anonymous):

i just do these percentage problems forward, and figure them out.

OpenStudy (anonymous):

what do you mean? how i understand is that he paid 22.40 for a sweatshirt that was discounted 30%. Is that correct?

OpenStudy (anonymous):

Yes..

OpenStudy (anonymous):

well then he pays for 70% of the regular price, which is $22.40

OpenStudy (anonymous):

correct?

OpenStudy (anonymous):

yep. you're doing it correctly.

OpenStudy (anonymous):

so then, the price needs to be 32 dollars because a 30% discount of 32 dollars yields 22.40. What was wrong with my reasoning?

hero (hero):

This is a simple problem of finding the number if you you it's percentage.. You could re-write the question as 22.70 is 70% of what? so is/of =%/100 ....22.70/x = 70/100 Cross Multiply: x = 32.43 So anhhuylex was right... Good job